Grade 10Mechanics

Find the radius of gyration of a uniform meter rule suspended from 0.1m oscillating as a compound pendulum. Mass=0.05kg,l=4.2*10^-2kgm^2.hence find the period of oscillation of the compound pendulum

Ukobo obiageli
9 Years agoGrade 10

1 Answer

Mercy Akintimehin
4 Years ago
Using the formula: Moment of inertia, I=mk^2 +mh^2(Where k is the radius of gyration)Hence, by collecting like termsK=√(I-mh^2÷m)K=√(4.2×10^-2 - 0.05×0.1^2 ÷0.05)K=2.88mPeriod of oscillation, T=2π√(I/mghT=2×22/7√(4.2×10^-2/0.05×9.8×0.1)T=5.8secsThanks 😊